7. Find the specific gravity, weight density, and volume of the unknown object if it



weighs 40 Ib in air and 30 Ib in water.



2. A 10,000 kg vehicle will round the having a radius of 150 meter. The curved road is



banked at 25 degrees with the horizontal, find the minimum coefficient of friction.



4. Find the acceleration and the centripetal force of a 14 kN car that rounds the curved



of radius of 50 meters having a speed of 8 km/h.



5. Two perfectly elastic balls, weighing 27 N and 18 N, approach each other with a



speed of 6 m/s and 11 m/s, respectively. What is the speed of the balls after the



collision?




1
Expert's answer
2021-11-16T13:50:11-0500

7. "40\ lb=178\ N" ; "30\ lb=133\ N"


Specific gravity "SG=\rho_{object}/\rho_{water}=178/133=1.34"


Density "\rho=1.34\cdot1000=1340\ (kg/m^3)"


"V=178/(1340\cdot9.8)=0.014\ (m^3)"



2. Suppose that "v=5\ (m/s)" . So, we have


"\mu=\frac{-v^2\cos25°+gR\sin25°}{v^2\sin25°+gR\cos25°}=\frac{-5^2\cdot\cos25°+9.8\cdot 150\cdot\sin25°}{5^2\cdot\sin25°+9.8\cdot150\cdot \cos25°}=0.45"



4. "a=v^2/r=2.22^2/50\approx0.1\ (m/s^2)" and "F=ma=(14000/9.8)\cdot0.1=141\ (N)"


5. "v_1=\frac{m_1-m_2}{m_1+m_2}u_1+\frac{2m_2}{m_1+m_2}u_2=\frac{2.8-1.8}{2.8+1.8}\cdot 6+\frac{2\cdot 1.8}{2.8+1.8}\cdot 11="


"=9.91\ (m/s)"

"v_2=\frac{2m_1}{m_1+m_2}u_1+\frac{m_2-m_1}{m_1+m_2}u_2=\frac{2\cdot 2.8}{2.8+1.8}\cdot 6+\frac{1.8-2.8}{2.8+1.8}\cdot 11="


"=4.91\ (m/s)"